Using Limits to Calculate a Derivative
Use the definition of the derivative to find f ′(x) if
f(x)= 1/(x-1)

Answer by fractalier(6550)

The derivative of
f(x)= 1/(x-1)
is
f(x) = -1/(x-1)^2
You use the limit formula
f'(x) = lim(h-->0) of (f(x+h) - f(x)) / h
to get there.